# Qubit-Efficient Quantum Annealing for Stochastic Unit Commitment

### v1

@@ III-B2 Equivalent QUBO Form and Hamiltonian
In the QPHR-ALM framework, the new QUBO form of the objective function is obtained in the previous iteration, so using a QA processor, the $\ell$-th iteration eigenstate $\left|\boldsymbol{x}^{\ell}\right\rangle$, where $\boldsymbol{x}^{\ell}$ is a binary qubit string, can then be measured. It is worth highlighting that a bit string is not the same as an eigenstate. For example, in a two-qubit system, when the measured result $\left|\boldsymbol{x}^{\ell}\right\rangle$ is $\left|00\right\rangle$, the bit string $\boldsymbol{x}^{\ell}$ is $00$, which means that the two binary variables $x_{1}^{\ell}$ and $x_{2}^{\ell}$ are 00. However, the eigenstate is $\left|00\right\rangle=\left|0\right\rangle\otimes\left|0\right\rangle=\begin{bmatrix}1&0&0&0\end{bmatrix}^{\mathrm{T}}$, which is not equivalent to $\boldsymbol{x}^{\ell}$. Therefore, to obtain the quantum state solution in the next iteration, we must consider the structure of the Hamiltonian. Due to the addition of an augmented penalty term, the Hamiltonian consists of two parts, an objective Hamiltonian $\boldsymbol{H}_{Obj}$ representing an objective function $f(\boldsymbol{x})$, and the PHR-based augmented Lagrangian Hamiltonian $\boldsymbol{H}_{PHR}$ standing for the PHR augmented penalty function.
$\boldsymbol{H}=\boldsymbol{H}_{Obj}+\boldsymbol{H}_{PHR}=\boldsymbol{H}_{Obj}+\sum_{i=1}^{w}\boldsymbol{H}_{PHR,i}$ (15)
@@
It is evident that $\boldsymbol{H}$ is not constant in each iteration. When an optimization problem is well-defined, it results in a clear expression of $f(\boldsymbol{x})$, ensuring the invariance of $\boldsymbol{H}_{Obj}$. However, $\boldsymbol{H}_{PHR}$ is not fixed, and it represents the aggregate of the Hamiltonians corresponding to all unsatisfied constraints. Additionally, since the $\max$ function in (13) cannot be directly embedded in QA-QPUs, $\boldsymbol{x}^{\ell}$ is employed to classify the PHR augmented penalty function from (12) into two cases, determining the expression for $\boldsymbol{H}_{PHR}$ for all inequality constraints.
@@
  • Case 1: Substituting the bit string $\boldsymbol{x}^{\ell}$ into $\lambda_{i}+\sigma g_{i}(\boldsymbol{x})$, if the result is less than or equal to 0, the bit string of the eigenstate $\left|\boldsymbol{x}^{\ell}\right\rangle$ satisfies the $i$-th inequality constraint. Then, according to (12), the relevant $\boldsymbol{H}_{PHR,i}$ is a diagonal matrix with component $-\left(\lambda_{i}{}^{2}/2\sigma\right)$. Actually, in this case, the PHR augmented Lagrangian Hamiltonian is equivalent to a constant term, so it does not affect the optimal solution but only influences the value of objective function. Therefore, it can be neglected directly.
@@
  • Case 2: When ${\lambda_{i}+\sigma g_{i}(\boldsymbol{x}^{\ell})>0}$, it implies that $\left|\boldsymbol{x}^{\ell}\right\rangle$ violates the current constraint, and the eigenstate result lies outside the feasible region. In this case, it is necessary to apply a related augmented penalty Hamiltonian to guide the next eigenstate sequence $\left|\boldsymbol{x}^{\ell+1}\right\rangle$ closer to the feasible region. Consequently, QUBO-formed expression of $\boldsymbol{H}_{PHR,i}$ can be obtained by (12).
@@
Consequently, when $\left|\boldsymbol{x}^{\ell}\right\rangle$ is known, the QUBO form of the PHR augmented Lagrangian function can be derived using (12) and (13). Next, convert the QUBO form into the Hamiltonian $\boldsymbol{H}_{PHR}$ by (5) and (6). After updating the new Hamiltonian, we can use it to evolve a new quantum state $\left|\boldsymbol{x}^{\ell+1}\right\rangle$. If the acquired quantum state satisfies the stop criterion (14c) or reaches the maximum iteration limit, the iteration can be terminated. Otherwise, proceed to adjust Lagrange multipliers and penalty coefficients relying on (14a) and (14b).
To sum up, by eliminating the slack variables, the qubit overhead of the QPHR-ALM algorithm is reduced from $NT+J+k\mathcal{F}$ to $NT+J$, which is independent to the number of Bender iterations.
@@
0:  LBO-formed $f(\boldsymbol{x})$ and linear binary constraints
0:  Optimal quantum state $\left|\boldsymbol{x}^{*}\right\rangle$
@@
3:     Obtain $\left|\boldsymbol{x}^{\ell}\right\rangle$ using QA algorithms
@@
5:        if $\lambda_{i}+\sigma g_{i}\left(\boldsymbol{x}^{\ell}\right)>0$ then
@@ III-C Quantum-based PHR-ADMM Algorithm
Although the QPHR-ALM method removes slack variables, reducing the qubit overhead to $NT+J$, this is still insufficient to meet the requirements of the actual SUC problem. For example, when scheduling 24-hour operation of a power system with 12 generator units hourly, the master problem, after applying BD, would involve at least 288 integer variables. Considering the discretized LBs, it might require a total of over 300 qubits after binay encoding. For large-scale power systems, the qubit requirement could escalate to thousands or even tens of thousands, far exceeding the capacity of current QPUs.
Since it is impractical to rely on a single quantum computer for the entire computational workload, the ADMM algorithm can be implemented for the SUC master problem. This approach decomposes a large-scale LBO-formatted integer programming master problem into several smaller sub-problems that can be efficiently managed by smaller scale QPUs, and the solution to the large global problem is obtained by coordinating the solutions to all the sub-problems.
For multi-unit, multi-period SUC problems, the D-ADMM approach can decompose the master problem by units, which means that the states for all time periods of each unit are grouped into the same block. Considering that the number of decomposed subproblems will exceed two, the multi-block D-ADMM algorithm for the master problem can be compactly expressed as
$\min$ $\sum_{m=1}^{M}f_{m}\left(\boldsymbol{x}_{m}\right)$ (16)
s.t. $\sum_{m=1}^{M}g_{i,m}\left(\boldsymbol{x}_{m}\right)\leq 0\text{ for }i=1,\ldots,w$
Based on a similar derivation process in Section III-B, we can easily obtain the PHR-based augmented Lagrangian function for the D-ADMM form after eliminating the slack variables.
$\!\!\!\!\!\!\mathcal{L}_{D}\!=\frac{1}{2\sigma}\!\sum_{i=1}^{w}\left(\!\left[\max\left\{\sigma\!\sum_{m=1}^{M}g_{i,m}\left(\boldsymbol{x}_{m}\right)\!+\lambda_{i},0\right\}\right]^{2}\!\!-\!\lambda_{i}{}^{2}\!\right)\\ +\sum_{m=1}^{M}f_{m}\left(\boldsymbol{x}_{m}\right)$ (17)
@@
$\boldsymbol{x}_{m}^{\ell+1}=$ $\arg\min_{\boldsymbol{x}_{m}^{\ell}}\mathcal{L}_{D}(\boldsymbol{x}_{1}^{\ell+1},\ldots,\boldsymbol{x}_{m-1}^{\ell+1},\boldsymbol{x}_{m}^{\ell},\boldsymbol{x}_{m+1}^{\ell},\ldots,\boldsymbol{x}_{M}^{\ell},$
@@
$\lambda_{i}^{\ell+1}=$ $\max\left\{\lambda_{i}^{\ell}+\sigma^{\ell}\sum_{m=1}^{M}g_{i,m}\left(\boldsymbol{x}_{m}^{\ell+1}\right),0\right\}\!\text{ for }i=1,\ldots,w$ (18b)
$\sigma^{\ell+1}=$ $\eta\sigma^{\ell}$ (18c)
$\mathcal{R}^{\ell}\!=\!$ $\left[\sum_{i=1}^{w}\!\left(\max\left\{\!-\frac{\lambda_{i}^{\ell}}{\sigma^{\ell}},\sum_{m=1}^{M}g_{i,m}\left(\boldsymbol{x}_{m}^{\ell+1}\right)\right\}\right)^{2}\right]^{1/2}\!\leq\delta$ (18d)
The convergence of QPHR-ADMM is based on D-ADMM. Its convergence properties remain as an open problem [36]. However, to achieve optimal convergence, the following criteria should ideally be met: 1) Strong convexity. 2) The penalty coefficient $\sigma$ should be selected within a suitable range. 3) The step size factor $\eta$ should be sufficiently small [36, 37].
Notably, as each block in (18a) can be serially solved on QPUs, the total QPU qubit overhead further reduces to $\max\{T,J\}$ as a constant.
Remark 1
Alternatively, the binary variables can be partitioned into $T$ blocks, resulting in qubits number $\max\{N,J\}$. As $T$ is conventionally fixed as 24 in power system operation, our setting aligns with the worst possible qubits number when $N\gg T$.
To sum up, the overall saving on qubits number is summarized in Table I, which is independent of the number of generators and the number of BD iterations.
TABLE I: Summary on the qubit number and computational paradigm of the algorithm under BD framework.
Method Theoretical Qubit Number Computation Paradigm
Basic QA $NT+J+k\mathcal{F}$ Quantum
QPHR-ALM $NT+J$ Quantum
QPHR-ADMM $\max\{T,J\}$ Quantum
III-D Framework of BD and QPHR-ADMM for SUC
Our proposed QPHR-ADMM algorithm, powered by PHR Augmented Lagrangian and quantum ADMM, is summarized in Fig. 2. The algorithm begins by decoupling the SUC problem using BD. This decoupling serves two primary purposes. First, it ensures that all scenarios within the second stage are independent and continuous in the sub-problems, which can be efficiently solved by allocating dedicated CPU to each scenario in parallel. Second, the first-stage master problem is decomposed by ADMM, allowing for sequential processing of each block on a single quantum processor. Each block is then solved by the PHR Augmented Lagrangian method without occupying extra qubits for slack variables introduced by inequality constraints.
The binary unit decisions from QPHR-ADMM iterations are then integrated into subproblems to optimize power outputs and update dual variables. The calculation of subproblems contributes to tightening the UB of BD. The process repeats until the BD bounds converge within a small tolerance.
@@ IV Simulations
This section validates the reliability of the algorithm through two optimization simulation cases based on QPHR-ADMM, and then verifies its feasibility and effectiveness for SUC.
IV-A Binary Integer Programming Example

### v2

@@ III-B2 Equivalent QUBO Form and Hamiltonian
In the QPHR-ALM framework, the new QUBO form of the objective function is obtained in the previous iteration. Therefore, the $\ell$-th iteration eigenstate $\left|\bm{x}^{\ell}\right\rangle$, where $\bm{x}^{\ell}$ is a binary qubit string, can then be measured using a QA-QPU. As shown by Fig. 1, to obtain the quantum state solution in the next iteration, we must consider the structure of the Hamiltonian. Due to the addition of an augmented penalty term, the Hamiltonian consists of two parts, an objective Hamiltonian $\bm{H}_{Obj}$ representing an objective function $f(\bm{x})$, and the PHR-based augmented Lagrangian Hamiltonian $\bm{H}_{PHR}$ standing for the PHR augmented penalty function.
$\bm{H}=\bm{H}_{Obj}+\bm{H}_{PHR}=\bm{H}_{Obj}+\sum_{i=1}^{w}\bm{H}_{PHR,i}$ (17)
@@
It is evident that $\bm{H}$ is not constant in each iteration. When an optimization problem is well-defined, it results in a clear expression of $f(\bm{x})$, ensuring the invariance of $\bm{H}_{Obj}$. However, $\bm{H}_{PHR}$ is not fixed, and it represents the aggregate of the Hamiltonians corresponding to all unsatisfied constraints. Additionally, since the $\max(\cdot)$ function in (15) cannot be directly embedded in QA-QPUs, $\bm{x}^{\ell}$ is employed to classify the PHR augmented penalty function from (14) into two cases, determining the expression for $\bm{H}_{PHR}$ for all inequality constraints:
@@
  • Case 1: Substituting the bit string $\bm{x}^{\ell}$ into $\lambda_{i}+\sigma g_{i}(\bm{x})$, if the result is less than or equal to 0, the bit string of the eigenstate $\left|\bm{x}^{\ell}\right\rangle$ satisfies the $i$-th inequality constraint. Then, according to (14), the relevant $\bm{H}_{PHR,i}$ is a diagonal matrix with component $\left(\lambda_{i}{}^{2}/2\sigma\right)$. Consequently, the PHR augmented Lagrangian Hamiltonian is equivalent to a constant term, so it does not affect the optimal solution but only influences the value of objective function. Therefore, it can be neglected directly.
@@
  • Case 2: When $\lambda_{i}+\sigma g_{i}(\bm{x}^{\ell})>0$}, it implies that $\left|\bm{x}^{\ell}\right\rangle$ violates the current constraint, and the eigenstate result lies outside the feasible region. In this case, it is necessary to apply a related augmented penalty Hamiltonian to guide the next eigenstate sequence $\left|\bm{x}^{\ell+1}\right\rangle$ closer to the feasible region. Consequently, QUBO-formed expression of $\bm{H}_{PHR,i}$ can be obtained by (14).
@@
Consequently, when $\left|\bm{x}^{\ell}\right\rangle$ is known, the QUBO form of the PHR augmented Lagrangian function can be derived using (14) and (15). Next, convert the QUBO form into the Hamiltonian $\bm{H}_{PHR}$ by (5) and (6). After updating the new Hamiltonian, we can use it to evolve a new quantum state $\left|\bm{x}^{\ell+1}\right\rangle$. If the acquired quantum state satisfies the stop criterion (16c) or reaches the maximum iteration limit, the iteration can be terminated. Otherwise, proceed to adjust Lagrange multipliers and penalty coefficients relying on (16a) and (16b).
In terms of qubit efficiency, as no slack variables are introduced for inequality constraints, the qubit overhead of the QPHR-ALM algorithm is reduced from $\underline{NT+J+\mathcal{M}\mathcal{P}+k\mathcal{F}}$ to $\underline{NT+J}$, which is independent of the increasing number of Benders’ cuts during iteration and the minimum up/down time constraints. The pseudo-code is provided in Algorithm 1.
@@
0: LBO-formed $f(\bm{x})$ and linear binary constraints
0: Optimal quantum state $\left|\bm{x}^{*}\right\rangle$
@@
3:  Obtain $\left|\bm{x}^{\ell}\right\rangle$ using QA algorithms
@@
5:   if $\lambda_{i}+\sigma g_{i}\left(\bm{x}^{\ell}\right)>0$ then
@@ III-C Quantum-based PHR-ADMM Algorithm
Although the QPHR-ALM method eliminates slack variables and reduce the qubit overhead to $NT+J$, it remains inadequate for practical SUC instances. For example, when scheduling a 24-hour operation on IEEE bus-118 system with 54 generator units hourly, the master problem, after applying BD, would involve at least 1296 integer variables. Since it is impractical to rely on a single quantum computer for the entire computational workload, the ADMM algorithm is adopted for solving SUC master problem. This approach decomposes a large-scale LBO-formatted integer programming master problem into several smaller sub-problems that can be efficiently managed by smaller scale QPUs, and the solution to the large global problem is obtained by coordinating the solutions to all the sub-problems.
In particular, for multi-unit, multi-period SUC, the D-ADMM approach decomposes the master problem either by generator units or time horizon. Under unit-wise decomposition, the states for all time periods of each unit are grouped into the same block, followed by a block accouting for the binary encoded continuous variable $\Upsilon^{k}$ with $J$ binary variables. As $T$ is conventionally fixed as 24 in power system operation, this setting aligns with the worst possible qubits number when $N\gg T$. Moreover, the ADMM block size must respect hardware embedding limits to avoid large minor embeddings that can increase chain breaks; for the current D-Wave Advantage QPU, each subproblem is preferably restricted to roughly 30–32 binary variables [49]. From this perspective, decomposing by generator units is the more reasonable and hardware-aware choice.
The multi-block D-ADMM algorithm for the compact master problem (10) can be expressed as
$\min$ $\sum_{m=1}^{M}f_{m}\left(\bm{x}_{m}\right)$ (18)
s.t. $\sum_{m=1}^{M}g_{i,m}\left(\bm{x}_{m}\right)\leq 0\text{ for }i=1,\ldots,w$
where $M$ is the number of ADMM blocks. Based on a similar derivation process in Section III-B, we can easily obtain the PHR-based augmented Lagrangian function for the D-ADMM form after eliminating the slack variables as,
$\mathcal{L}_{D}\!=\frac{1}{2\sigma}\!\sum_{i=1}^{w}\left(\!\left[\max\left\{\sigma\!\sum_{m=1}^{M}g_{i,m}\left(\bm{x}_{m}\right)\!+\lambda_{i},0\right\}\right]^{2}\!\!-\!\lambda_{i}{}^{2}\!\right)\\ +\sum_{m=1}^{M}f_{m}\left(\bm{x}_{m}\right)$ (19)
@@
$\bm{x}_{m}^{\ell+1}=$ $\arg\min_{\bm{x}_{m}^{\ell}}\mathcal{L}_{D}(\bm{x}_{1}^{\ell+1},\ldots,\bm{x}_{m-1}^{\ell+1},\bm{x}_{m}^{\ell},\bm{x}_{m+1}^{\ell},\ldots,\bm{x}_{M}^{\ell},$
@@
$\lambda_{i}^{\ell+1}=$ $\max\left\{\lambda_{i}^{\ell}+\sigma^{\ell}\sum_{m=1}^{M}g_{i,m}\left(\bm{x}_{m}^{\ell+1}\right),0\right\}\!\text{ for }i=1,\ldots,w$ (20b)
$\sigma^{\ell+1}=$ $\begin{cases}\sigma^{\ell},&\text{if }\mathcal{R}^{\ell+1}<\varrho\mathcal{R}^{\ell},\\[2.0pt] \eta\,\sigma^{\ell},&\text{otherwise}\end{cases}$ (20c)
$\mathcal{R}^{\ell}\!=\!$ $\left[\sum_{i=1}^{w}\!\!\left(\max\!\left\{\!-\frac{\lambda_{i}^{\ell}}{\sigma^{\ell}},\!\sum_{m=1}^{M}\!g_{i,m}\!\left(\bm{x}_{m}^{\ell+1}\right)\!\right\}\!\right)^{\!2}\right]^{\!1/2}\!\!\leq\!\delta$ (20d)
Notably, as each block in (20a) can be serially solved on QPUs, the total QPU qubit overhead further reduces from $NT+J$ to $\underline{\max\{T,J\}}$ as a constant.
III-D Summary on the BD and QPHR-ADMM Framework
Our proposed QPHR-ADMM algorithm, powered by PHR Augmented Lagrangian and quantum ADMM, is summarized in Fig. 2 and the overall saving on qubits numbers is summarized in Table II, which is independent of the number of generators and the number of BD iterations. The algorithm begins by decoupling the SUC problem using BD. This decoupling serves two primary purposes for both CPU and QPU efficiency. First, it ensures that all scenarios within the second stage are independent and continuous in the sub-problems, which can be efficiently solved by allocating dedicated CPUs to each scenario in parallel. Second, the first-stage master problem is decomposed by ADMM, allowing for sequential processing of each block on a single quantum processor. Each block is then solved by the PHR Augmented Lagrangian method without occupying extra qubits for slack variables introduced by Benders’ cuts. The binary unit decisions from QPHR-ADMM iterations are then integrated into subproblems to optimize power outputs and update dual variables. The calculation of subproblems contributes to tightening the UB of BD. The process repeats until the BD bounds converge within a small tolerance.
TABLE II: Summary on the qubit number of QPHR-ADMM under BD framework.
Method Qubit Number
Basic QA $NT+J+\mathcal{M}\mathcal{P}+k\mathcal{F}$
QPHR-ALM $NT+J$
QPHR-ADMM $\max\{T,J\}$
@@ III-D Summary on the BD and QPHR-ADMM Framework
III-E Discussion on the Algorithm Convergence
The convergence of the proposed algorithm depends on convergence properties of its individual components, inclduing BD, QA solver, augmented Lagrangian, and D-ADMM. First, BD guarantees finite termination and global optimality for mixed-integer programs under standard assumptions, notably exact subproblem solutions and the generation of valid cuts [15]. For the QA solver on the master problem, convergence guarantees are primarily available in idealized adiabatic regimes. When the annealing process is sufficiently slow and appropriate regularity conditions hold, the final state remains close to the ground state, with potentially exponential error decay under stronger smoothness assumptions[1]. In our setting, the Benders master problem in this work is discrete and non-strongly convex; recent results [46] show that, for block-structured integer programs, an augmented Lagrangian formulation with a sufficiently large but finite penalty parameter can yield an exact relaxation and valid lower bounds, provided that the subproblems are solved accurately. At last, for convex problem, the convergence of ADMM typically requires the penalty $\sigma$ to be chosen within a suitable range and the step size factor $\eta$ to be sufficiently small [18, 17]. By contract, ADMM lacks general convergence guarantees in nonconvex or discrete settings and is therefore used here as a coordination mechanism rather than as a provably convergent solver. Nevertheless, empirical evidence in power system applications suggests that stable performance can be achieved by initializing $\sigma$ conservatively and increasing it gradually [50, 27].
@@ IV Simulations
This section validates the reliability of the proposed algorithm through synthetic binary problems as well as SUCs based on a 4-generator power system demonstrator and the IEEE bus-118 system. The proposed and baseline QC algorithms are tested on the QPU solver, D-Wave Advantage_system 4.1. We also use the Gurobi solver on a classical computer with an Intel Core i7-12700H processor as reference. To start, we verify the convergence of the proposed QPHR-ADMM on a synthetic MIP instance in Appendix -A, which also demonstrates that QA yields more accurate solutions than QAOA in this setting.
IV-A 4 Generators Scenario-based SUC Example
The power system considered in this section consists of four generators, an equivalent wind generator, and a load. The output of the wind turbine is influenced by wind speed over a 24-hour period, which follows a Weibull probability distribution. Additionally, the load demand is uncertain and is assumed to follow a Beta distribution[22].
IV-A1 Theoretical Analysis on Qubit Overhead
Considering that the time horizon of the SUC is $24$ hours, the master problem after BD involves $24\times 4=96$ decision variables for the total commitment statuses. Additionally, $J=12$ and $\mathcal{F}=13$ are set to allocate qubits for discretizing the LB and slack variables, with the corresponding precision coefficients specified as $0.004$ and $0.005$, respectively. Therefore, If the minimum up/down time constraints are temporarily ignored, a total of $108+13k$ qubits are required in the basic QA where $k$ is the number of Bender’s iterations. In contrast, since no binary-encoded slack variables are introduced, the number of qubits required for QPHR-ALM can be determined as $108$. In contrast, based on the QPHR-ADMM algorithm, we decompose the master problem into five unit-based blocks. The first four blocks contain the on//off decision variables for each generator over the $24$-hour period, while the last block is used to compute the discretized LB value. As a result, these five blocks require $24$, $24$, $24$, $24$, and $12$ qubits. As the D-ADMM framework can be implemented in serial, it implies that the QPHR-ADMM algorithm requires only a QPU with a capacity of $24$ qubits to finish tasks that would otherwise demand $108+13k$ qubits.
If the minimum up- and down-time requirements are set to $T_{g}^{U}=[6,4,3,2]$ and $T_{g}^{D}=[6,4,3,2]$ for 4 units, respectively, the four generators will give rise to a total of 162 inequality constraints. If the same slack-variable setting that $\mathcal{P}=13$ is applied to each constraint, the resulting binary slack variables would require 2106 qubits in total, which is far beyond the capability of the Basic QA. In contrast, QPHR-ADMM still requires only 24 qubits.
IV-A2 Convergence and Sensitivity Analysis
Fig. 3 shows that the trends in the LB values and the corresponding errors obtained using the QPHR-ADMM method (with different parameter $\sigma^{0}$) and Basic QA during the first iteration of BD for the $100$ sampled scenarios. It should be emphasized that the results in Fig. 3 are obtained without minimum up- and down-time constraints. Including these constraints increases the qubit requirement beyond the minor-embedding capability of current quantum annealing hardware, making Basic QA infeasible.
Figure 3: Stopping criteria error (top) and BD lower bound (bottom) versus ADMM iterations within the first BD iteration for $\sigma^{0}\in\{0.0015,0.0067,0.0130\}$. Solid lines: D-Wave; dotted lines: Gurobi; purple: Basic QA ($\sigma^{0}=0.0015$); red line: true value.
QPHR-ADMM demonstrates satisfactory convergence performance. For example, as shown in Fig. 3, the LB value gradually narrows down to the optimal value, and the stopping criteria error eventually approaches zero during the QPHR-ADMM iterations. It indicates that the master problem will converge to the correct solution. Additionally, This figure also demonstrates the impact of the initial penalty parameters on convergence performance. For instance, a larger $\sigma^{0}$ can reduce iterations, thereby saving computational time. However, due to the uncertain convergence of QPHR-ADMM, setting a smaller $\sigma^{0}$ is necessary to ensure more reliable convergence, as discussed before. Moreover, the measured results from basic QA often fail to converge to the correct ground-state solution, leading to erroneous LB values as shown by the purple curves. Such issues occur due to the high qubit demand, which leads to longer embedding chains. The increased chain length reduces the effective chain strength of certain qubits, resulting in broken chains in current D-Wave QA-QPUs [37]. Consequently, employing QPHR-ADMM not only lowers qubit overhead per annealing process but also plays a critical role in maintaining solution accuracy, even under limited qubit resources.
Table III illustrates the convergence of upper and lower bounds in the BD method under different numbers of scenarios, sets of constraints, and wind penetration level. The UB results are obtained by aggregating all the subproblems, which are processed exclusively by classical computers. The LB can be computed using either QPUs or CPUs for reference. As shown in the table, after introducing time-coupled start-up/down constraints and higher wind power penetration, the proposed method is still able to converge. Note that there exists a slight discrepancy between the LBs computed by Gurobi and the proposed QPHR-ADMM, due to the precision level of the discretized LB being set at $0.004$.
At last, Fig. 4 presents the sensitivity of convergence time on parameters $\sigma^{0}$ and $\eta$ in QPHR-ADMM with minimum up/down constraints. As outlined in Section III-E, smaller penalty coefficients $\sigma^{0}$ and step size factors $\eta$ tend to promote more stable convergence, whereas increasing them can substantially reduce run time and accelerate convergence; however, excessively large values may result in non-convergence. Therefore, we recommend selecting the parameters within the red box.
TABLE III: The BD Results for SUC Problems with 4 Generators.
Assumptions: No minimum up/down time constraints
$10$ scenarios $100$ scenarios
0 $27.856$ 0 0 $27.644$ 0 0
11 $11.501$ $11.056$ $11.056$ $11.290$ $10.844$ $10.844$
22 $11.501$ $11.501$ $11.500$ $11.290$ $11.290$ $11.288$
Assumptions: With $T_{g}^{U}=[6,4,3,2]$ and $T_{g}^{D}=[6,4,3,2]$
$10$ scenarios $100$ scenarios
0 $27.856$ 0 0 $27.644$ 0 0
11 $11.896$ $11.367$ $11.368$ $11.665$ $11.136$ $11.136$
22 $12.201$ $11.957$ $11.956$ $11.823$ $11.698$ $11.700$
33 $12.058$ $11.924$ $11.924$ $11.665$ $11.665$ $11.664$
44 $12.040$ $11.929$ $11.928$
55 $11.896$ $11.896$ $11.896$
Assumptions: With $T_{g}^{U}=[6,4,3,2]$, $T_{g}^{D}=[6,4,3,2]$ and $1.1P^{Wind}$
$10$ scenarios $100$ scenarios
0 $27.660$ 0 0 $27.442$ 0 0
11 $11.761$ $11.170$ $11.172$ $11.532$ $10.933$ $10.932$
22 $12.054$ $11.822$ $11.824$ $11.681$ $11.562$ $11.560$
33 $11.915$ $11.790$ $11.788$ $11.532$ $11.532$ $11.532$
44 $11.901$ $11.794$ $11.796$
55 $11.761$ $11.761$ $11.760$
Figure 4: The stability map for 4 generator UC problem with 100 scenarios (Gray blocks mean non-convergence)
Figure 5: Average direct QPU access time cost for different QUBO sizes (equivalent to the qubit requirement of a single ADMM block).
IV-B Large-scale SUC Problem
This section evaluates the performance of QPHR-ADMM on large-scale SUC problems using IEEE bus-118 system with 54 generating units. For modeling simplicity, the minimum up/down time of all generating units are assumed to be 4 hours. Table IV summarizes the optimization results of the SUC problem using classical and quantum computing methods. The problem contains 1296 binary decision variables, and the basic QA is infeasible due to hardware limitations. However, the QPHR-ADMM enables solving the problem with only 24 qubits. In terms of solution quality, the classical BD method using Gurobi achieves a total cost of 1728.88 k\$. The proposed QPHR-ADMM attains an average cost of 1734.26 k\$, with a standard deviation of 1.12 k\$ over multiple runs. This corresponds to a cost increase of approximately 0.31% compared with the classical BD solution. The observed gap mainly results from occasional QA errors in unit commitment decisions, which lead to suboptimal on/off statuses in a small number of time periods. These results demonstrate that QPHR-ADMM maintains convergence behavior over large system and delivers near-optimal solutions while significantly reducing the quantum resource requirement.
TABLE IV: Cost Comparison Between Classical and Quantum Method.
Classcial Computing Quantum Computing
Cost (Gurobi) Cost (Gurobi + BD) Binary Variables Average Cost $\pm$ std (QPHR-ADMM + BD) Qubit Overhead
$1728.80\,k\mathdollar$ $1728.88\,k\mathdollar$ 1296 $1734.26\pm 1.12\,k\mathdollar$ 24
IV-C Discussion and Prospects for Quantum Advantage
In the context of QC, its primary potential advantage lies in accelerating computations for power system scheduling. This section discusses such prospects under both current hardware constraints and anticipated future developments. Fig. 5 illustrates the solving time of D-Wave QA processors for QUBO problems of varying sizes, considering only direct QPU access time and ignoring solution errors in larger-scale annealing.
For the smaller-scale QUBO problems, classical computers exhibit a clear advantage in computational speed over quantum processors. However, as the problem size increases, the computational time for classical methods escalates significantly due to the NP-hard nature of integer programming. In contrast, the measured QPU access time of QA shows relatively weak sensitivity to problem size growth. These empirical observations suggest that, beyond a certain problem scale (e.g., on the order of $100$ equivalent qubits), quantum annealing may become competitive with classical approaches in terms of solver execution time, provided that solution quality is maintained.
From a system-level perspective, it should be emphasized that the QPU access time reported in Fig. 5 does not represent the end-to-end computational cost of QA. In practical deployments, additional time overheads are introduced by classical–quantum communication, minor-embedding construction and queueing. In contrast, classical solvers such as Gurobi can be deployed locally with negligible communication latency. As a result, under current hardware conditions, the practical wall-clock time of QA-based solutions generally exceeds the pure QPU execution time and does not yet outperform state-of-the-art classical solvers.
Beyond these timing-related considerations, current hardware limitations also impose constraints on algorithmic implementation. In particular, due to chain-break risks arising from minor embedding on D-Wave QPUs, the block size adopted in QPHR-ADMM must presently remain relatively small to ensure solution stability. These constraints are primarily attributable to limited qubit connectivity, rather than to inherent limitations of the algorithmic framework. With continued advances in quantum annealing hardware, such as increased connectivity, reduced chain-break rates, and larger numbers of reliable logical qubits, the same framework is expected to support substantially larger block sizes, and enabling more effective utilization of quantum acceleration in large-scale power system optimization.
V Conclusion
This paper investigates the feasibility of solving scenario-based SUC through quantum computation. Building on the classic BD framework, we propose a novel QPHR-ADMM algorithm that integrates the principles of QPHR-ALM and D-ADMM. The proposed approach enables independent acceleration of the master problem on a quantum annealer, while the subproblems are solved in parallel on CPUs. Additionally, the algorithm mitigates potential quantum solver failures and reduces qubit overhead, maintaining it at a constant level rather than scaling linearly with the number of Benders iterations, generators, and additional constraints. By demonstrating the applicability of QA to address synthetic and SUC problems, this study establishes a foundational framework for addressing complex mixed-integer inequality constrained optimizations in power system operation. Furthermore, the framework can be extended to other BD-based power system optimization problems with large-scale binary master variables, such as robust UC or transmission expansion planning, although problem-specific modeling and convergence challenges may require further investigation.
-A Binary Integer Programming Example